# On the new and accurate (Goudsmit-Saunderson) model for describing e-/e+ multiple Coulomb scattering

### v1

@@ 2.3.3 More than one scattering case
$\frac{\tilde{w}^{2}}{0.5(s/\lambda)+2}=\left\{\begin{array}[]{ll}1.347+t(0.209364-t(0.45525-t(0.50142-t0.081234)))&\quad\mathrm{if}\;\lambda<10\\ &\\ -2.77164+t(2.94874-t(0.1535754-t0.00552888))&\quad\mathrm{otherwise}\\ \end{array}\right.$ (45)
@@ 2.3.4 Screening parameter
with $137.03561/\alpha=$ 137.0356, $r_{0}=2.81794\times 10^{-13}$ [cm], $u=1.660538\times 10^{-24}$ [g] and $C_{TF}$ as given by Eq.(48) was used and $\beta^{2}$ was neglected. According to Eq.(50), $A^{(M)}$ can be expressed by using this material dependent $b_{c}^{*}$ parameter as
@@
where $\chi_{\alpha}$ is Moliere’s screening angle (the screening parameter is $A^{(M)}=\chi_{\alpha}^{2}/4$) of the compound that builds up from $N_{e}$ elements, $\chi_{\alpha_{i}}$ are the screening angles characterising the $i-th$ components with corresponding atomic number $Z_{i}$ and $n_{i}$ is the proportion by number of the $i-th$ atom ($n_{i}=\rho\mathcal{N}_{\mathcal{A}}/M_{i}=\rho/(uA_{M_{i}})$ (Note, that $Z^{2}$ is replaced by $Z(Z+\xi_{e})$ with $\xi_{e}\in[0,1]$ to account scattering from sub-threshold inelastic events). The individual Moliere’s screening angles are expressed in terms of the corresponding Thomas-Fermi screening angle $\chi_{0_{i}}$ as
@@
(where the $1/\beta^{2}$ part was dropped from the last term similarly to Eq.(53)). The Thomas-Fermi screening angle for the $i-th$ components with atomic number $Z_{i}$ is
@@
where $n_{i}$ is the fraction of the $i-th$ atom in the compound by number $n_{i}=$ number-of-atoms-per-unit-volumei per total number of atoms per volume. (Note, that only the ratios are important since a possible constant term will be cancelled later: (i) $Z_{X}/Z_{S}$ and $Z_{E}/Z_{S}$ will appear in Z-s; (ii) $b_{c},\chi_{cc}^{2}$ will contain a $Z_{S}/A$ term where $A=\sum_{i}n_{i}A_{i}$ relative molecular mass with $A_{i}$ being the molar mass of the $i-th$ atom).
@@ 2.4.2 Pre-computation of the $q^{2+}(u)$ transformed PDF
As mentioned above, sampling tables are generated over an appropriate 2D grid of discrete $s/\lambda,G_{1}s/\lambda$ values. Practical considerations in [1] suggested the $s/\lambda\in[1,10^{5}]$ and $G_{1}s/\lambda\in[0.001,0.99]$ limits with $64$ linearly spaced points on $\ln{(s/\lambda)}$ scale and $15$ $G_{1}s/\lambda$ values spaced linearly.
@@ 2.4.3 Run-time: sampling MSC deflections from GS angular distributions
In the third, i.e. at least two scattering, case the pre-computed sampling tables can be utilised to generate samples of the transformed variable $u$ according to the corresponding $q^{2+}(u;s,\lambda,G_{1},\tilde{a})$ transformed PDF. The final $\mu(u)$ sample can then be computed by the inverse transform given by Eq.(36).
@@ 3.1 Energy loss correction to the GS angular distributions
were derived in [2] for the case of utilising the screened Rutherford DCS for computing the GS angular distributions (as our case) with $\tilde{E}=E_{0}-\Delta E/2$ mid-point energy along the step, $\epsilon=\Delta E/E_{0}$ fractional energy loss along the step and $\tilde{\tau}=\tilde{E}/m_{e}$ is the mid-point energy in electron rest-mass $m_{e}$ units. Interestingly, only $E_{\text{eff}}$ has dependence on the stopping power $L(E)$ through $b(\tilde{E})$ but even that can be neglected as $b(\tilde{E})$ is proportional to $\mathrm{d}C(E)/\mathrm{d}E$ with $C(E)=L(E)\beta^{2}$ being only a slowly varying function with $E$. Introducing then $\tilde{\epsilon}=\Delta E/\tilde{E}$ one can write the above as222there seems to be two typos in the EGSnrc implementation: one for $E_{\text{eff}}$ having $(\tilde{\tau}^{2}+2\tilde{\tau}+3)$ instead of $(\tilde{\tau}^{2}+3\tilde{\tau}+2)$ and the second for $s_{\text{eff}}$ having $\tilde{\epsilon}^{2}$ instead of $\tilde{\epsilon}\epsilon$ though none of them is significant
$\begin{split}E_{\text{eff}}&=\tilde{E}\left[1-\frac{\tilde{\epsilon}^{2}}{24}\frac{(5\tilde{\tau}^{2}+10\tilde{\tau}+6)}{(\tilde{\tau}^{2}+3\tilde{\tau}+2)}\right]\\ s_{\text{eff}}&=s\left[1-\frac{\epsilon\tilde{\epsilon}}{6}\frac{(\tilde{\tau}^{4}+4\tilde{\tau}^{3}+7\tilde{\tau}^{2}+6\tilde{\tau}+4)}{(\tilde{\tau}+1)^{2}(\tilde{\tau}+2)^{2}}\right]\end{split}$ (77)
@@ 3.2 Accounting spin-effect
Figure 2: The $R_{\text{MR}}$ Mott-to-Rutherford elastic DCS ratio, that shows the effect of the spin to the DCS, for $Au$ and $Al$ target atoms, electron and position projectiles at various projectile kinetic energy. See APPENDIX A.6 for the details of the computation.
@@
This assumes that spin and screening effects are decoupled which is motivated by the fact that $R_{MR}(\theta)$ goes to unity (i.e. no spin effects) at small scattering angles (especially with decreasing projectile energies), while screening affects most significantly these small angle scatterings (especially at lower projectile energies) [22, 21]. Furthermore, spin effects are more significant at larger target atomic numbers in case of electron projectile as demonstrated in Fig. 2. Therefore, Eq.(80), that was also adopted in EGSnrc to account the combined spin and screening effects, might be a reasonable approximation at projectile energies $>~{}100$ [keV] in case of low-Z while requires $>~{}1$ [MeV] for high-Z materials [3].
@@
More might be seen regarding the applicability of the above factorisation by comparing the DCS-s computed with the different approximations summarised in Table 1 and shown in Fig. 4 for low- ($Al$) and high-Z ($Au$) target atoms at varying projectile kinetic energies. In general, one can see that the simple Rutherford (R) and screened Rutherford (SR) DCS-s are the same at large scattering angles while they differ only at small $\theta$ values where the effect of the screening is clearly visible. The Rutherford (R) and Mott (M) DCS are the same at these small scattering angles (and diverge in the same way due to the lack of screening) while they differ at larger $\theta$ values where spin effect can be important as expected from Fig. 2. The spin-corrected screened Rutherford (or screening corrected Mott) DCS (SR×$\times\text{R}_{\text{MR}}$) agrees with the Mott (M) DCS at large scattering angles while follows the screened Rutherford (SR) one at smaller $\theta$ values accounting both the spin and and screening effects respectively.
@@
All the four DCS-s agree well in case of low-Z ($Al$) target at smaller projectile energies as the spin effect to the DCS tends to vanish for low-Z targets with decreasing kinetic energies (see Fig. 2 right). At larger projectile energies, when spin effects starts to be more pronounced even for low-Z targets, the SR×$\times\text{R}_{\text{MR}}$ DCS keeps following very well the ELSEPA one thanks to the spin correction. In contrast to the low-Z targets, spin effects might be important in case of high-Z targets also at lower projectile energies (see Fig. 2 left) leading to large differences between those not including (e.g. R, SR) spin effects and the others. Furthermore, a relatively large projectile kinetic energy $\sim 128-1024$ [keV] is needed for reaching a good agreement between the SR×$\times\text{R}_{\text{MR}}$ (same for M) and ELSEPA DCS-s in case of high-Z targets even at these large scattering angles.
@@
A simple exponential screening is used to account the screening of the potential of the nucleus by the presence of the atomic electrons when deriving the screened Rutherford (SR) DCS. The limitation of this approximation is responsible for the relatively large differences observable at small scattering angles between the SR×$\times\text{R}_{\text{MR}}$ and ELSEPA DCS-s that later includes accurate screening computed from the Dirac-Fock model of the atom. This difference reduces more rapidly with increasing projectile energies in case of low-Z targets and practically vanishes at $\sim 128$ [keV] for scattering angles $\theta>\sim 1^{\circ}$, while this requires larger $\sim 1024$ [keV] projectile energy for high-Z targets.
Therefore, the SR×$\times\text{R}_{\text{MR}}$ DCS given by Eq.(80), requires larger projectile kinetic energies $\sim 1$ [MeV] to be a reasonable approximation for high-Z target materials while somewhat lower $\sim 100$ [keV] energies are already sufficient for low-Z targets.
@@ 3.3 Screening parameter correction
The accuracy of the corresponding multiple scattering angular distributions can be increased even further though. To this end, it is important to recognise that the multiple scattering angular distributions are strongly determined by the first transport cross section 555according to Eq.(A.23) the first moment of the GS angular distribution, i.e. the mean scattering angle along an $s$ path is $<\cos(\theta)>_{GS}=\exp(-sG_{\ell=1}/\lambda)=\exp(-s/\lambda_{1})$ where $\lambda_{1}$ is the first transport mean free path that is $1/\lambda_{1}=\mathcal{N}\sigma_{1}$ with the number of scattering centres (atoms) per unit volume $\mathcal{N}$ and first transport cross section $\sigma_{1}$ as given by Eq.(10) thus the large angle behaviour of the DCS  [9]. It is obvious from Eq.(10), that the small scattering angle part of the DCS do not contribute significantly to the first transport cross section due to the $[1-\cos(\theta)]$ factor. Therefore, the differences between the SR×$\times\text{R}_{\text{MR}}$ and the ELSEPA DCS-s, observable at small $\theta<\sim 1^{\circ}$ scattering angles and discussed above, have only mild effect to the corresponding multiple scattering angular distributions. Furthermore, the $\theta>\sim 1^{\circ}$ accuracy of the SR×$\times\text{R}_{\text{MR}}$ DCS, therefore the accuracy of the corresponding first transport cross section and the related multiple scattering angular distributions, can be increased further by applying a $\kappa$ correction to Moliere’s screening parameter $A^{(M)}$ (used so far in $\mathrm{d}\sigma^{(SR)}/\mathrm{d}\Omega$ and given by Eq.(68)) such that
@@
holds for the SR×$\times\text{R}_{\text{MR}}$ DCS with the corrected $A=\kappa A^{(M)}$ screening parameter. In other words, the SR×$\times\text{R}_{\text{MR}}$ DCS with the corrected screening parameter $A=\kappa A^{(M)}$ reproduces the first transport cross section computed from the corresponding accurate ELSEPA DCS. A data base of the $\kappa$ correction factor has been generated both for $e^{-}$ and $e^{+}$ projectiles, target atomic numbers Z = $1-98$ over an appropriate kinetic energy grid as suggested in [3] 666an $E_{\text{kin}}$ kinetic energy grid on $1-100$ [keV] with 16 log-spacing points where linear interpolation on $\ln(E_{\text{kin}})$ is used and a second $\beta^{2}$ grid above from $E_{\text{kin}}=100$ [keV] ($\beta^{2}=0.300546$) to $\beta^{2}=0.9999$ ($E_{\text{kin}}=50.5889$ [MeV]) with 16 linear-spacing points where linear interpolation on $\beta^{2}$ is used. The numerical ELSEPA DCS and $R_{MR}$ Mott-to-Rutherford ratio were computed for $e^{-}/e^{+}$ at each target atomic number, projectile energy point and the corresponding $\kappa(Z,E_{\text{kin}},e^{-}/e^{+})$ correction factors were determined by solving the above Eq.(81) numerically. Some examples of the $\kappa=A/A^{(M)}$ corrections factors are shown in Fig.3 for different target atoms and projectile energies. The correction factors deviates significantly from unity only at high-Z targets and lower projectile energies where the larger angle behaviour of the spin-corrected SR×$\times\text{R}_{\text{MR}}$ DCS deviates from the corresponding ELSEPA DCS as discussed in the previous section. These precomputed, element-wise correction factors $\kappa(Z,E_{\text{kin}},e^{-}/e^{+})$ are stored in separate files for $e^{-}/e^{+}$ over the pre-defined kinetic energy grid and used at initialisation for computing the material (compound) dependent screening correction factors for all materials present in the geometry using Eq.(70) (separate factors for $e^{-}/e^{+}$ over the pre-defined kinetic energy grid).
Figure 3: Some $\kappa=A/A^{(M)}$ screening correction factors calculated by solving Eq.(81) numerically for the given projectile (left $e^{-}$; right $e^{+}$), target atom and kinetic energy to obtain the corrected $\kappa A^{(M)}$ screening parameter that satisfies Eq.(81) when used in the SR part of the SR×$\times\text{R}_{\text{MR}}$ DCS. Similar correction is used in EGSnrc (see Figure 13. in [3] or Figure 1. in [23]) and the reported correction factors are in a good agreement with the presented values (though different PWA DCS were used in [3, 23] as reference which explains the small observable differences).
It must be noted, that it is the SR×$\times\text{R}_{\text{MR}}$ DCS that gives the same first transport cross section with the corrected screening parameter as the corresponding ELSEPA DCS and not the pure SR DCS. Therefore, the pure SR DCS based (analytical) expressions for the elastic mean free path ($\lambda$ as given by Eq.(66)) and first transport coefficients ($G_{\ell=1}$ as given by Eq.(A.40), used to compute the first transport mean free path as defined by Eq.(66), do not provide a result that would agree with the corresponding more accurate ELSEPA DCS based value even when using the corrected screening parameter. On the same time, the accuracy of the first (and second) transport mean free path is important as it determines the first (and second) GS moment (see Eqs.(A.23,A.24)). These are especially important in the accurate electron-step algorithm discussed in Section 4 as the first- and second-order spatial moments of the step are strongly determined by $\xi\equiv sG_{\ell=1}/\lambda$ and $\gamma\equiv sG_{\ell=2}/\lambda$ (without energy loss along the $s$ step but see more at Section 4).
@@
Therefore, element-wise correction factors to the first transport cross sections (and to the second-to-first) have also been pre-computed based on the relation between the above mentioned analytical SR (using the corrected screening) and the more accurate numerical ELSEPA DCS based values. The corresponding element-wise correction factors are also stored in separate files for $e^{-}/e^{+}$ over the pre-defined kinetic energy grid, loaded at initialisation and utilised to prepare multiplicative correction factors to the SR model based analytical values such that the more accurate ELSEPA DCS related results are reproduced. Therefore, this material, projectile kinetic energy and type (i.e. $e^{-}$ or $e^{+}$) dependent multiplicative correction factors can be utilised at run time (in combination with the corrected screening parameter) to obtain more accurate $\xi\equiv sG_{\ell=1}/\lambda$ and $\gamma\equiv sG_{\ell=2}/\lambda$ values for the electron-step computation similarly to EGSnrc [3] .
@@
Figure 4: Differential Cross Sections (DCS-s) for Coulomb scattering of electrons on $Au$ and $Al$ target atoms at different projectile kinetic energies. See Table 1 and the text.
@@ 3.4 Scattering power correction
TO BE COMPLETED…
It has been discussed at the end of Section 2.3.4, that $Z(Z+1)$ must be replaced with $Z(Z+\xi_{e})$ and an appropriate $\xi_{e}\in[0,1]$ value must be chosen such that it accounts only the angular deflections from sub-threshold delta ray productions that are not simulated explicitly. This is to avoid double-counting as the $\xi_{e}=1$ value includes the contributions to the angular deflection also from the explicitly generated delta electron production. so called scattering power correction will be introduced [4].

### v2

@@ 2.3.3 More than one scattering case
[…]
@@ 2.3.4 Screening parameter
with $1/\alpha=137.0356$, $r_{0}=2.81794\times 10^{-13}$ [cm], $u=1.660538\times 10^{-24}$ [g] and $C_{TF}$ as given by Eq.(48) was used and $\beta^{2}$ was neglected. According to Eq.(50), $A^{(M)}$ can be expressed by using this material dependent $b_{c}^{*}$ parameter as
@@
where $\chi_{\alpha}$ is Moliere’s screening angle (the screening parameter is $A^{(M)}=\chi_{\alpha}^{2}/4$) of the compound that builds up from $N_{e}$ elements, $\chi_{\alpha_{i}}$ are the screening angles characterising the ithi-th components with corresponding atomic number $Z_{i}$ and $n_{i}$ is the proportion by number of the ithi-th atom ($n_{i}=\rho\mathcal{N}_{\mathcal{A}}/M_{i}=\rho/(uA_{M_{i}})$ (Note, that $Z^{2}$ is replaced by $Z(Z+\xi_{e})$ with $\xi_{e}\in[0,1]$ to account scattering from sub-threshold inelastic events). The individual Moliere’s screening angles are expressed in terms of the corresponding Thomas-Fermi screening angle $\chi_{0_{i}}$ as
@@
(where the $1/\beta^{2}$ part was dropped from the last term similarly to Eq.(53)). The Thomas-Fermi screening angle for the ithi-th components with atomic number $Z_{i}$ is
@@
where $n_{i}$ is the fraction of the ithi-th atom in the compound by number $n_{i}=$ number-of-atoms-per-unit-volumei per total number of atoms per volume. (Note, that only the ratios are important since a possible constant term will be cancelled later: (i) $Z_{X}/Z_{S}$ and $Z_{E}/Z_{S}$ will appear in Z-s; (ii) $b_{c},\chi_{cc}^{2}$ will contain a $Z_{S}/A$ term where $A=\sum_{i}n_{i}A_{i}$ relative molecular mass with $A_{i}$ being the molar mass of the ithi-th atom).
@@ 2.4.2 Pre-computation of the $q^{2+}(u)$ transformed PDF
As mentioned above, sampling tables are generated over an appropriate 2D grid of discrete $s/\lambda,G_{1}s/\lambda$ values. Practical considerations in [1] suggested the $s/\lambda\in[1,10^{5}]$ and $G_{1}s/\lambda\in[0.001,0.99]$ limits with $64$ linearly spaced points on $\ln{(s/\lambda)}$ scale and 1515 $G_{1}s/\lambda$ values spaced linearly.
@@ 2.4.3 Run-time: sampling MSC deflections from GS angular distributions
[…]
@@ 3.1 Energy loss correction to the GS angular distributions
were derived in [2] for the case of utilising the screened Rutherford DCS for computing the GS angular distributions (as our case) with $\tilde{E}=E_{0}-\Delta E/2$ mid-point energy along the step, $\epsilon=\Delta E/E_{0}$ fractional energy loss along the step and $\tilde{\tau}=\tilde{E}/m_{e}$ is the mid-point energy in electron rest-mass $m_{e}$ units. Interestingly, only $E_{\text{eff}}$ has dependence on the stopping power $L(E)$ through b(E~)b($\tilde${E}) but even that can be neglected as b(E~)b($\tilde${E}) is proportional to $\mathrm{d}C(E)/\mathrm{d}E$ with $C(E)=L(E)\beta^{2}$ being only a slowly varying function with $E$. Introducing then $\tilde{\epsilon}=\Delta E/\tilde{E}$ one can write the above as (keeping the same $\mathcal{O}(\epsilon^{2})$ precision as above also when moving to $\tilde{\epsilon}$) 222there seems to be a typo in the EGSnrc implementation ($E_{\text{eff}}$ having $(\tilde{\tau}^{2}+2\tilde{\tau}+3)$ instead of $(\tilde{\tau}^{2}+3\tilde{\tau}+2)$) though not causing significant difference
$\begin{split}E_{\text{eff}}&=\tilde{E}\left[1-\frac{\tilde{\epsilon}^{2}}{24}\frac{(5\tilde{\tau}^{2}+10\tilde{\tau}+6)}{(\tilde{\tau}^{2}+3\tilde{\tau}+2)}\right]\\ s_{\text{eff}}&=s\left[1-\frac{\tilde{\epsilon}^{2}}{6}\frac{(\tilde{\tau}^{4}+4\tilde{\tau}^{3}+7\tilde{\tau}^{2}+6\tilde{\tau}+4)}{(\tilde{\tau}+1)^{2}(\tilde{\tau}+2)^{2}}\right]\end{split}$ (77)
@@ 3.2 Accounting spin-effect
Figure 2: The $R_{\text{MR}}$ Mott-to-Rutherford elastic DCS ratio, that shows the effect of the spin to the DCS, for $Au$ and AlAl target atoms, electron and position projectiles at various projectile kinetic energy. See APPENDIX A.6 for the details of the computation.
@@
This assumes that spin and screening effects are decoupled which is motivated by the fact that $R_{MR}(\theta)$ goes to unity (i.e. no spin effects) at small scattering angles (especially with decreasing projectile energies), while screening affects most significantly these small angle scatterings (especially at lower projectile energies) [22, 21]. Furthermore, spin effects are more significant at larger target atomic numbers in case of electron projectile as demonstrated in Fig. 2. Therefore, Eq.(80), that was also adopted in EGSnrc to account the combined spin and screening effects, might be a reasonable approximation at projectile energies >100>~100 [keV] in case of low-Z while requires >1>~1 [MeV] for high-Z materials [3].
@@
[…]
@@
[…]
@@
[…]
@@ 3.3 Screening parameter correction
[…]
@@
[…]
@@
Therefore, element-wise correction factors to the first transport cross sections (and to the second-to-first) have also been pre-computed based on the relation between the above mentioned analytical SR (using the corrected screening) and the more accurate numerical ELSEPA DCS based values. The corresponding element-wise correction factors are also stored in separate files for $e^{-}/e^{+}$ over the pre-defined kinetic energy grid, loaded at initialisation and utilised to prepare multiplicative correction factors to the SR model based analytical values such that the more accurate ELSEPA DCS related results are reproduced. Therefore, this material, projectile kinetic energy and type (i.e. $e^{-}$ or $e^{+}$) dependent multiplicative correction factors can be utilised at run time (in combination with the corrected screening parameter) to obtain more accurate $Q_{1}\equiv sG_{\ell=1}/\lambda$ and $\gamma\equiv G_{\ell=2}/G_{\ell=1}$ values for the electron-step computation similarly to EGSnrc [3] .
@@
Figure 4: Differential Cross Sections (DCS-s) for Coulomb scattering of electrons on AuAu and AlAl target atoms at different projectile kinetic energies. See Table 1 and the text.
@@ 3.4 Scattering power correction
It has been discussed in Section 2.3.4, that $Z^{2}\to Z(Z+\xi)$ was introduced with the usual $\xi=1$value to incorporate the angular deflections due to interactions with atomic electrons or more precisely the angular deflections due to the sub-threshold ionisations along the simulation step that are not modelled explicitly. However, secondary electrons with initial energy above the production threshold are generated and the corresponding angular deflections of the primary particle are accounted. Therefore, the corresponding contribution should not be included in the description of elastic scattering to avoid double counting and only the angular deflections due to sub-threshold ionisations should be accounted. It means that an appropriate, production threshold dependent $0\leq\xi\leq 1$ value should be used. It will be assumed in the following that using $\xi=1$ is a valid approximation under the continuous slowing down case, i.e. when there is no delta ray production during the simulation777though the validity of this usual approximation is not clear.
Kawrakow investigated this issue and derived an appropriate correction in [4]. Exactly that correction is implemented thus summarised below (all details can be found in section IV of [4]). First he obtained expressions for the mass scattering power (the rate of increase of the mean square scattering angle with traversed mass thickness) due to multiple Coulomb and Moller scatterings, i.e. due to scattering on the atomic potential and on the atomic electrons. He obtained [4]
$\frac{T_{R}(\zeta)}{\rho}=\frac{4\pi N_{A}r_{e}^{2}}{\beta^{2}\tau(\tau+2)}\,\frac{Z(Z+\zeta)}{A}\,g_{R}(\chi_{a})$
with
$g_{R}(A^{(M)})=\left(1+2A^{(M)}\right)\ln\left(1+\frac{1}{A^{(M)}}\right)-2.$
for the Coulomb scattering part of the mass scattering power, using the SR DCS with $A^{(M)}$ screening parameter (given in Section 2.3.4). Depending on the value of $\xi$ it might include some, all or no contributions from electron-electron scattering. He obtained for the electron-electron part, that corresponds to the explicitly produced secondary electrons (i.e. above threshold), [4]
$\frac{T_{M}(\tau_{c})}{\rho}=\frac{4\pi N_{A}r_{e}^{2}}{\beta^{2}\tau(\tau+2)}\,\frac{Z}{A}\,g_{M}(\tau,\tau_{c}),$
with
$g_{M}(\tau,\tau_{c})=\ln\frac{\tau}{2\tau_{c}}+\left[\frac{(\tau+2)}{(\tau+1)^{2}}\right]^{2}\ln\frac{2(\tau-\tau_{c}+2)}{\tau+4}-\left[\frac{(\tau+2)^{2}}{4}+\frac{(\tau+2)(\tau+1/2)}{(\tau+1)^{2}}\right]$
$\times\ln\frac{(\tau+4)(\tau-\tau_{c})}{\tau(\tau-\tau_{c}+2)}+\frac{(\tau-2\tau_{c})(\tau+2)}{2}\left[\frac{1}{\tau-\tau_{c}}-\frac{1}{(\tau+1)^{2}}\right].$
using the Møller scattering and its free electron at rest approximation. $\tau,\tau_{c},N_{A},r_{e},Z,A$ are the primary electron kinetic energy in rest mass units, the secondary electron production kinetic energy threshold in rest mass units, Avogadro’s constant, electron classical radius, target atomic number and relative atomic mass.
Assuming then that an elastic mass scattering power with the original $\xi=\xi_{0}=1$ parameter $T_{R}(\xi_{0})/\rho$ correctly describes the total contributions to the scattering power (i.e. both the real Coulomb and the total (above and below) electron-electron contributions), the elastic scattering part should have a parameter $\xi_{e}$ that satisfies the following equation
$\frac{T_{R}(\xi_{0})}{\rho}=\frac{T_{M}(\tau,\tau_{c})}{\rho}+\frac{T_{R}(\xi_{e})}{\rho}$
This states that the real Coulomb plus the total electron-electron contribution ($T_{R}(\xi_{0})/\rho$) should be equal to the above threshold electron-electron ($T_{M}(\tau,\tau_{c})/\rho$) plus an elastic scattering contribution that includes both the real Coulomb and the sub-threshold electron-electron contributions ($T_{R}(\xi_{e})/\rho$). As we would like to have exactly this latter, we can solve the above equation for $\xi_{e}$ that gives
$\xi_{e}=\xi_{0}-\frac{g_{M}(\tau,\tau_{c})}{g_{R}(A^{(M)})}$
As $g_{M}(\tau,\tau_{c})$ is 0 whenever $\tau$ is below the secondary production threshold (i.e. $\tau\leq\tau_{c}$) then $\xi_{e}=\xi_{0}$ with the usual $\xi=1$ value covering all electron-electron interaction contribution while otherwise $g_{M}(\tau,\tau_{c})/g_{R}(A^{(M)})\in(0,1]$ leading to $\xi_{e}<\xi_{0}$ removing the above threshold contributions.
This $\xi$ parameter is used in $Z_{S}$ defined in Eq.(65) that transforms now with this scattering power correction(spc) to
$Z_{S}^{\text{spc}}=\sum_{i=1}^{N_{e}}n_{i}Z_{i}(Z_{i}+\xi_{e})=\sum_{i=1}^{N_{e}}n_{i}Z_{i}\left[Z_{i}+\xi_{0}-\frac{g_{M}(\tau,\tau_{c})}{g_{R}(A^{(M)})}\right]=Z_{S}\left[1-\frac{\sum_{i=1}^{N_{e}}n_{i}Z_{i}}{Z_{S}}\frac{g_{M}(\tau,\tau_{c})}{g_{R}(A^{(M)})}\right]$
with $Z_{S}=\sum_{i=1}^{N_{e}}n_{i}Z_{i}(Z_{i}+\xi_{0})$. This multiplicative, kinetic energy and production threshold dependent scattering power correction parameter to $Z_{S}$ is computed at initialisation for each material-cuts couple above an appropriate discrete kinetic energy threshold (from $2E_{c}$ for electron and $E_{c}$ for positron) and interpolated at run time. Note, that only the elastic mean free path depends on $Z_{S}$ through Moliere’s $b_{c}$ parameter as given by Eq.(66) and Eq.(67) ($Z_{S}$ from $\chi_{cc}^{2}$ and $b_{c}$ cancel each other in the case of $A^{(M)}$; see Eq.(68)). Therefore, we apply this scattering power correction only to the elastic mean free path at run time after calculated by using Eq.(66).
It has been mentioned above (and also discussed by Kawrakow in [4]) that the full validity of the $Z^{2}\to Z(Z+\xi)$ replacement to incorporate angular deflections from electron-electron interactions to the elastic scattering is not clear. However, the main advantage of the above scattering power correction to avoid double counting is that it makes the simulation of angular deflections independent from the secondary electron production threshold used in the ionisation. This can be especially important in case of light (low Z) materials as the contribution from sub-threshold electron-electron interactions to the total angular deflection can be significant compared to the Coulomb scattering. Moreover, it must be noted that neglecting those angular deflections can lead to inaccurate modelling results independently from the applied, i.e. multiple or single, Coulomb scattering model.